A bond offers a $50 coupon, has a face value of $1,000, and 15 years until maturity. If the interest rate is 4.5%, what is the value of the bond?
Answer #1

Value of bond=Annual coupon*Present value of annuity factor(4.5%,15)+$1000*Present value of discounting factor(4.5%,15)

=50*10.73954573+$1000*0.516720442
=$1053.70(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=50[1-(1.045)^-15]/0.045

=50*10.73954573

2.Present value of discounting factor=1000/1.045^15

=$1000*0.516720442